OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)

OpenStudy (anonymous):

slope of lin =y2-y1/x2-x1 =-3-3/1+5 =-6/6 =-1 equation of line is y=mx+c m=slope y=-1x+c as line passes thru (1,-3) -3=-1+c c=-2 y=-1x-2 to get x intercept keep y=0 -x-2=0 x=-2 point is(-2,0)

OpenStudy (anonymous):

do u got it give me medal if u like that sol yar